22=6(2x-3)-3(5x-8)
how do I find what x is?

Answer by longjonsilver(2297)

22=6(2x-3)-3(5x-8)
22 = 12x - 18 - 15x + 24
22 = -3x - 18 + 24
22 = -3x + 6
22 + 3x = 6
3x = -16
x = -16/3

jon.